OpenStudy (anonymous):

A dance floor has a perimeter of 60 meters. The dance floor is 20 meters long. Use this information to help you answer questions 1-5. Please explain your answers with words, and show your work, including computations. 1. Write an equation using variables to represent the perimeter of the dance floor. 2. Solve the equation to find the dimensions of the dance floor.

OpenStudy (anonymous):

So we have P = 60m, L = 20m. Guess this is a rectangle. The equation for the perimeter P of a rectangle is \[P = 2L + 2W\], where L = length, and W = width. Here, we allready know P and L => we have the equation\[60 = 2*20 + 2*x\]. Now, we will solve for x, which is the width of the rectangle and we obtain\[x = (60-40)/2 = 20/2 = 10\], which means that the dimensions of the dance floor is L*W = 20*10.

OpenStudy (anonymous):

thnx!!!!!!! :D ;D